Find the area of the smaller loop r=2-4sin0, symmetrical with Oy.


1
Expert's answer
2021-04-14T14:57:50-0400

We know,


Area can be evaluated as "\int rd \theta"


We have given that "r = 2-4sin\theta"


Hence "I = \int_{0}^{2\pi} (2-4sin\theta) d\theta"


"I = [_{0}^{2\pi} (2\theta+4cos\theta)"


We get "I = 4\pi + 4 - 0 - 4"


Hence, "I = 4\pi"